# ExPath: Targeted Pathway Inference for Biological Knowledge Bases via Graph Learning and Explanation

### v1

@@ F.1. Classifier ablation
Table 5 presents the results of PathExplainer when the classifier model is changed. In our PathMamba, the fidelity+ score is reduced to less than one-tenth, indicating that the extracted subgraph alone produces nearly identical results. This suggests that ExPath is capable of extracting biologically functional pathways sufficiently. Our PathMamba exhibits a relatively low fidelity+, which may imply that it has strong representational power, allowing it to learn meaningful representations even from subgraphs that were not extracted.
Appendix G Case Study
Figure 6. Comparison of subgraphs extracted from the TCR signaling pathway using two different methods. The TCR Subgraph on the left is from the RSS method, and the TCR Subgraph on the right is from the proposed method. The subgraph nodes and their signaling modules are colored in red. The disruptions within signaling paths are marked in green boxes.
G.1. T cell receptor (TCR) signaling pathway
The T cell receptor (TCR) signaling pathway is a cornerstone of adaptive immunity, orchestrating antigen-specific T cell activation, clonal expansion, and effector differentiation (Huse, 2009). This pathway is initiated upon engagement of the TCR complex with peptide-MHC ligands, triggering a cascade of intracellular signaling events mediated by the Src-family kinases LCK and FYN, leading to phosphorylation of the immunoreceptor tyrosine-based activation motifs (ITAMs) within the CD3 and $\zeta$-chain subunits (Gaud et al., 2018). Subsequent recruitment and activation of ZAP-70 further amplify downstream signaling through the LAT signalosome, engaging multiple adaptor proteins and second messengers that regulate key pathways, including calcium mobilization, Ras-MAPK, and NF-$\kappa$B signaling, which collectively drives gene transcription, metabolic reprogramming, and cytoskeletal remodeling necessary for T cell function (Shah et al., 2021).
@@ G.1. T cell receptor (TCR) signaling pathway
Precise modulation of these signaling cascades is critical for maintaining immune homeostasis, as dysregulation is implicated in a spectrum of immune disorders, including autoimmunity, primary immunodeficiencies, and T cell malignancies, where aberrant activation or attenuation of TCR signaling disrupts immune tolerance, facilitates chronic inflammation, or drives oncogenic transformation. Analysis of the molecular intricacies of TCR signaling helps therapeutic interventions, including immune checkpoint modulation, CAR-T (Rafiq et al., 2020) cell engineering, and small-molecule inhibitors aimed at restoring immune balance and targeting immune-related diseases.
G.2. Results and discuss
@@
TCR Subgraph A: The RSS Method. As shown in Figure 6, the subgraph on the left, generated by the RSS method, distributes high scores uniformly across a broad range of nodes within the TCR pathway. However, this hints at unnatural, fragmented signal propagation, as highlighted by numerous discrete red-marked nodes. The absence of coherent signaling continuity, as indicated by the disrupted green-boxed regions, suggests that the method fails to prioritize biologically meaningful regulatory modules. Since its broader coverage spans multiple branches of the pathway without emphasizing critical molecular hubs, limiting its utility in pinpointing key functional perturbations.
TCR Subgraph B: The Proposed Method. In contrast, as shown in Figure 6, the subgraph on the right extracted by our method exhibits a strong focus on the PI3K-AKT signaling axis (Vara et al., 2004) and the downstream components of the NF-$\kappa$B (Dolcet et al., 2005) pathway, as highlighted by the coherent red-marked path. These regions are believed to be crucial for regulating T cell survival, proliferation, and cytokine production (Chen et al., 2013). Notably, the subgraph includes key regulatory genes such as MAPK1, MAP3K8, and NFKB1 (D’Souza et al., 2008; Cartwright et al., 2016) within a compact set of prioritized nodes. The enrichment of these molecular hubs suggests that our method effectively captures biologically significant signaling elements, aligning with well-established immune regulatory mechanisms.
In our case study, the proposed method provides a focused selection of key regulatory pathways, emphasizing PI3K-AKT and NF-$\kappa$B signaling and their downstream effectors, which are crucial for modulating immune responses. In contrast, the RSS method, while providing broader pathway coverage, lacks specificity, making it less suitable for pathway analyses requiring mechanistic interpretation.

### v2

@@ G.2 Classifier evaluation
Table 7 presents the results of PathExplainer when the classifier model is changed. In our PathMamba, the fidelity- score is reduced to less than one-tenth, indicating that the extracted subgraph alone produces nearly identical results. The lower fidelity+ score also shows strong representational power of PathMamba, as it can still perform reasonably well even when important subgraphs are removed. For our objective of identifying function-specific pathways, subgraph sufficiency is particularly important, which our method achieves effectively.
Appendix H Case Study
H.1 T cell receptor (TCR) signaling pathway
[…]
@@ H.1 T cell receptor (TCR) signaling pathway
Precise modulation of these signaling cascades is critical for maintaining immune homeostasis, as dysregulation is implicated in a spectrum of immune disorders, including autoimmunity, primary immunodeficiencies, and T cell malignancies, where aberrant activation or attenuation of TCR signaling disrupts immune tolerance, facilitates chronic inflammation, or drives oncogenic transformation. Analysis of the molecular intricacies of TCR signaling helps therapeutic interventions, including immune checkpoint modulation, CAR-T (Rafiq, Hackett, and Brentjens 2020) cell engineering, and small-molecule inhibitors aimed at restoring immune balance and targeting immune-related diseases.
H.2 Results and Discussion
@@
TCR Subgraph A: The RSS Method. As shown in Figure 6, the subgraph on the left, generated by the RSS method, distributes high scores uniformly across a broad range of nodes within the TCR pathway. However, this hints at unnatural, fragmented signal propagation, as highlighted by numerous discrete red-marked nodes. The absence of coherent signaling continuity, as indicated by the disrupted green-boxed regions, suggests that the method fails to prioritize biologically meaningful regulatory modules. Since its broader coverage spans multiple branches of the pathway without emphasizing critical molecular hubs, it limits the utility in pinpointing key functional perturbations.
TCR Subgraph B: The Proposed Method.
[…]
Appendix I Implementation Details
We performed a grid search over key hyperparameters to ensure optimal performance. Table 8 summarizes the values considered for each parameter.
Table 8: Hyperparameter settings.
Parameter Value
Batch size 32
Epochs 50
Cross-validation folds 10
Learning rate 0.001
Weight decay 5e-4
Number of layers {1, 2}
Hidden size {32, 64, 128}
Walk length {4, 8, 16}